# ToM-SWE: User Mental Modeling for Software Engineering Agents

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled coding agents to perform complex software engineering tasks, from code generation (Jiang et al., 2024) and debugging (Tian et al., 2024) to system design (Kovacic et al., 2025) and optimization (Gao et al., 2024). However, despite their impressive technical capabilities, coding agents often struggle with a fundamental aspect of software development: effective communication and collaboration with human developers.
The core limitation is that current systems lack explicit mechanisms for modeling and predicting human intent in long-horizon, multi-turn interactions (Kim et al., 2023). Unlike human developers who naturally build mental models of their collaborators’ goals, preferences, and constraints through various tasks (Tomasello, 2009), coding agents lack the mechanism to infer and acquire the underlying user intentions from the surface-level, which in real-world interactions are often ambiguous, incomplete, or context-dependent (Levinson, 1983). Furthermore, current coding agents typically operate in a stateless manner, treating each session as independent rather than maintaining persistent context about the user’s evolving goals and conversation history. This paradigm often leads to wasted effort and misunderstandings, and in high-stakes settings can result in erroneous, or even unsafe outcomes.
Figure 1: An example of how the ToM-SWE framework can help the SWE agent to model the user’s mental state and provide more useful suggestions based on the previous interaction history. When facing a casual starting instruction, the SWE agent simply generates code yet fails to meet the users’ requirement for the task. However, with the ToM agent, the SWE agent can first consult the ToM agent that persists user mental state across multiple sessions, and then act more aligned with the user’s preferences and constraints.
@@
To evaluate ToM-SWE’s effectiveness, we introduce the Stateful SWE benchmark, the first benchmark that allows agents to leverage realistic conversation histories across multiple coding sessions to track user mental state. In this setting, agents interact with an LLM-powered user simulator and receive synthesized interaction histories to guide their reasoning. To stay consistent with prior evaluation, we borrow the original issues from SWE-bench (Chowdhury et al., 2024), reframing them as casual starting instructions paired with different user profiles and interaction histories. The agent needs to understand user intent, preferences and constraints either through interacting with the LLM-powered user simulator or inferring from the past interaction histories. Unlike existing benchmarks such as SWE-bench, which primarily assess technical problem-solving, Stateful SWE evaluates an agent’s ability to sustain meaningful interactions over time. While Ambiguous SWE-bench (Vijayvargiya et al., 2025) tests ambiguity resolution by underspecifying instructions, it does not capture long-term memory demands.
@@
We evaluate our ToM-SWE framework on both our newly introduced stateful SWE benchmark as well as the stateless Ambiguous SWE-bench (Vijayvargiya et al., 2025), building our agent (ToMCodeAct) using the OpenHands platform (Wang et al., 2025), an open-source framework for developing SWE agents. We show that ToMCodeAct outperforms the OpenHands SOTA CodeAct agent (Wang et al., 2024) on both benchmarks. On the ambiguous SWE-bench, ToMCodeAct agent achieves 63.4% issue resolved rate compared to CodeAct agent’s 51.9% (+11.5% improvement). On the stateful SWE-bench, ToMCodeAct agent achieves $57.4\%$ (+43.9%) task resolved rate compared to CodeAct agent’s $13.5\%$. Furthermore, ToMCodeAct agent achieves substantially better user satisfaction scores of 3.62 compared to CodeAct agent’s 2.57 (+41% improvement, automatically measured through user simulators that evaluate preference alignment, communication, etc.). The results highlight the importance of modeling user in real-world software development scenarios to respect specific user preferences and constraints beyond task completion.
@@
Finally, we conduct an in-the-wild human study with professional developers to validate the practical effectiveness of our approach and find that ToM agent’s suggestions are useful 86% of the time. Unlike previous work that conducts human studies on predefined tasks (Wu et al., 2025; Qian et al., 2025), we recruit 17 software developers to use the ToM-enhanced OpenHands CLI (Wang et al., 2025) for their everyday, self-chosen coding tasks over three weeks. Each time the SWE agent communicates with the ToM agent, the developers choose to accept, partially accept, or reject the ToM agent’s suggestions. Developers can also provide feedback through a public Slack channel anytime during the study. Besides the high acceptance rate, we learn from the developers’ feedback that ToM agent can often provide useful and valuable suggestions that make users workflow more efficient (e.g., “Please add pytest for the new function” or “keep your code edit minimal”). Furthermore, ToM agent can even provide novel and preference aligned suggestions (e.g., “refactor the code following Linux development philosophy”).
@@ 2 ToM-SWE: Pairing SWE Agent with ToM Agent
Consider existing setups of software engineering agents such as SWE-agent (Yang et al., 2024) and OpenHands CodeAct (Wang et al., 2024; 2025). At time step $t$ in coding session $i$, an agent receives an observation $o_{t}\in O$ either from the environment (e.g., terminal output, file contents, test results) or the user (e.g., user instructions, feedback). Then the agent takes an action $a_{t}\in A$ (e.g., code edits, shell commands) following some policy $\pi(a_{t}|c^{i}_{t})$, where $c^{i}_{t}=(o_{1},a_{1},\ldots,o_{t-1},a_{t-1},o_{t})$ is the context available to the agent in the coding session $i$ until time step $t$.
@@ 3 Stateful SWE Benchmark
Figure 3: Overview of the stateful SWE benchmark. We first collect profiles following the three steps outlined above. We then create instances by pairing the user simulator of different profiles with SWE-bench issues. Original SWE-bench issue descriptions are rephrased by the user simulator into casual, natural-sounding starting instructions. The agent solves the task under the same environment and tests of the original instance with access to the previous interaction histories with the user and could interact with the simulated user for extra information.
@@ 4.1 Experiment Setup
Our experiments involve three agent variants: (1) CodeAct agent: The baseline implementation following Wang et al. (2024), which operates without explicit user modeling capabilities. (2) RAGCodeAct agent: An enhanced version that encourages the coding agent to proactively retrieve relevant information from previous interaction history, serving as the single agent paradigm to manage both the coding task and user modeling task simultaneously. (3) TomCodeAct agent: Our proposed approach that pair coding agent with a theory-of-mind agent, maintaining explicit user mental models and adapting behavior accordingly. We evaluate all agents using three state-of-the-art language models: Claude Sonnet 4, Claude 3.7 Sonnet and Qwen3-480B (Qwen3). We use Claude Sonnet 4 for the theory-of-mind agent across all experiments and use BM25 to retrive relevant information for both the RAGCodeAct agent and the TomCodeAct agent (see Appendix A.7 for complete model specifications and hyperparameters).
@@
As described in Section 3, we evaluate our approach on two complementary benchmarks: the Ambiguous SWE benchmark and our Stateful SWE benchmark. Both benchmarks use GPT-5 powered simulated users and evaluate agents over 500 instances with a maximum of 100 interaction turns per task. Note that the user simulator has access to both the complete issue description and the hints for the issue from the original SWE-bench issues. Together, these benchmarks evaluate complementary scenarios: ambiguous formal specifications requiring clarification versus informal user instructions with available context. We report task resolved rates for both benchmarks and additionally report user simulator satisfaction scores for the Stateful SWE benchmark.
@@ 4.2 Benchmark Results
Figure 4: Agent performance comparison across models. TomCodeAct agent consistently outperforms CodeAct agent across both benchmarks and model variants, with the largest performance gap observed in the Stateful SWE benchmark using Claude Sonnet 4.
Figure 4 shows the resolved rates across all model-agent combinations. For example, TomCodeAct agent maintains its lead with 63.4% resolution rate using Claude Sonnet 4 versus CodeAct agent’s 51.9% on the Ambiguous SWE benchmark. And TomCodeAct agent achieves 57.4% resolution rate with Claude Sonnet 4 compared to CodeAct agent’s 13.5%, representing a 43.9 percentage point improvement on the Stateful SWE benchmark.
@@
Table 1 shows user satisfaction scores for the Stateful SWE benchmark. TomCodeAct agent achieves the highest satisfaction ratings across all models.
@@
Besides the success of the TomCodeAct agent, we additionally have the following findings: (1) Theory-of-mind reasoning over past user-agent interactions is essential for understanding current user intent while simply retrieving raw session data provides limited help. We observe that while RAGCodeAct agent also has access to the previous raw session data, it still underperforms the ToMCodeAct agent across all models. This is especially true when the base model for SWE agent is not powerful enough to handle both the coding task and the user modeling task simultaneously. With Claude 3.7 Sonnet, RAGCodeAct agent even hurts the performance of the SWE agent (task resolved rate drops from 18.7% to 14.4%).
(2) User satisfaction does not equal solving the task. From Table 1, we observe the mismatch between task resolution rate and user satisfaction between RAGCodeAct agent and CodeAct agent. To further investigate the relationship between task resolution rate and user satisfaction, we separate the user satisfaction scores into three categories: High (3.5-5), Medium (2-3.5) and Low (1-2). And we focus on the cases where resolving the task disagrees with the user satisfaction, i.e., failed to resolve the task but user satisfaction is high (F+H) and resolved the task but user satisfaction is medium and low (S+M and S+L). Figure 5 provides a detailed analysis of agent performance across different categories along with the example feedback from the user simulator in Table 2. We observe that the ToMCodeAct agent has the highest F+H rates, indicating that even if the task is not resolved, the ToMCodeAct agent can still achieve higher user satisfaction. It is also interesting to see that the RAGCodeAct agent has the highest S+M rates, suggesting that modeling user preferences done wrong can lead to worse user satisfaction even if the task is resolved.
Table 2: Examples of Task Resolution and User Satisfaction Mismatches
Category Reason User Simulator Example Feedback (SWE-bench issue ID)
Fail + High Meaningful progress “precise fix with minimal back-and-forth, aligned well with my preferences” (matplotlib__matplotlib-22865)
Good communication “Asked for all key details up front (version, minimal repro, affected APIs), matching the preferred workflow” (django__django-16256)
Success + Med Ignored user’s preferred tools “Ignored the preferred typing style… Did not provide a descriptive commit message” (sympy__sympy-22456)
Poor communication style “You didn’t provide a brief summary…” (scikit-learn__scikit-learn-13779)
Figure 5: Task resolution and user satisfaction disagreement. F+H: the proportion of cases where the user simulator statisfaction score is high given the task is not resolved. S+M: the proportion of cases where the user simulator statisfaction score is medium or low given the task is resolved.
@@ 4.3 Cost Efficiency of ToM Agent
For each ToM base model (GPT-5 nano, GPT-5 mini, GPT-5, Claude 3.7, Claude 4), we run the TomCodeAct agent on 100 sampled Stateful SWE-bench instances. We then report the resolved rate (%) alongside the average session cost (USD), which includes the full end-to-end cost of using the ToM agent during problem solving. As shown in Figure 6, even very efficient ToM base models substantially improve performance while adding only a small fraction of the overall session cost. For instance, GPT-5 nano reaches 38.0% at only \$0.02 per session. The inset compares the average SWE cost per session (\$1.08) with the ToM consultation cost when using Claude 4 (\$0.17), showing that ToM accounts for approximately 16% of the total. Practitioners can choose ToM base models that balances cost and quality for their budget.
Figure 6: Resolved rate vs avg. cost with ToM agents models solving per instance.
5 Online Human Study
To validate ToM-SWE’s effectiveness in real-world settings, we conduct a three-week human study with 17 software developers who regularly use the baseline OpenHands CLI. We enhanced their familiar CLI environment with our TomCodeActAgent (powered by Claude Sonnet 4) to evaluate real-world effectiveness. In total, we collected 209 coding sessions from participants freely working on their own real-world, day-to-day coding tasks, in which there were 174 instances where the ToM agent provided suggestions.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled coding agents to perform complex software engineering tasks, from code generation (Jiang et al., 2026) and debugging (Tian et al., 2024) to system design (Kovacic et al., 2025) and optimization (Gao et al., 2024). However, despite their impressive technical capabilities, coding agents often struggle with a fundamental aspect of software development: effective communication and collaboration with human developers.
The core limitation is that current systems lack explicit mechanisms for modeling and predicting human intent in long-horizon, multi-turn interactions (Kim et al., 2023; Zhou et al., 2024). Unlike human developers who naturally build mental models of their collaborators’ goals, preferences, and constraints through various tasks (Tomasello, 2009), coding agents lack the mechanism to infer and acquire the underlying user intentions from the surface-level instructions. Furthermore, current coding agents typically operate in a stateless manner, treating each session as independent rather than maintaining persistent context about the user’s evolving goals and conversation history. This paradigm often leads to wasted effort and misunderstandings, and in high-stakes settings can result in erroneous, or even unsafe outcomes.
Figure 1: An example of how the ToM-SWE framework can help the SWE agent to model the user’s mental state and provide more useful suggestions based on the previous interaction history. When facing the same starting instruction (“Make me a website…”), the SWE agent simply generates code yet fails to meet the users’ requirement for the task. However, with the ToM agent, the SWE agent can first consult the ToM agent that maintains a persistent model of the user mental state across multiple sessions, and then act more aligned with the user’s preferences and constraints.
@@
To enable evaluation of coding agents in real-world settings, where agents must adapt to different users over long-horizon, multi-session interactions to solve software engineering tasks. We therefore introduce the Stateful SWE benchmark, the first benchmark that allows agents to leverage realistic conversation histories across multiple coding sessions to track user mental state. In this setting, agents interact with an LLM-powered user simulator and receive synthesized interaction histories to guide their reasoning. To stay consistent with prior evaluation, we borrow the original issues from SWE-bench (Jimenez et al., 2024), reframing them as ambiguous starting instructions paired with different user profiles and interaction histories. The agent needs to understand user intent, preferences and constraints either through interacting with the LLM-powered user simulator or inferring from the past interaction histories. Unlike existing benchmarks such as SWE-bench, which primarily assess technical problem-solving, Stateful SWE evaluates an agent’s ability to sustain meaningful interactions over time.
@@
We evaluate our ToM-SWE framework on both our newly introduced stateful SWE benchmark as well as the stateless Ambiguous SWE-bench (Vijayvargiya et al., 2026). We build our agent (ToMCodeAct) using the OpenHands platform (Wang et al., 2025), an open-source framework for developing SWE agents. We show that ToMCodeAct outperforms the OpenHands SOTA CodeAct agent (Wang et al., 2024) on both benchmarks. On the ambiguous SWE-bench, ToMCodeAct agent achieves 63.4% issue resolved rate compared to CodeAct agent’s 51.9% (+11.5% improvement). On the stateful SWE-bench, ToMCodeAct agent achieves $57.4\%$ (+43.9% task resolved rate compared to CodeAct agent’s $13.5\%$. Furthermore, ToMCodeAct agent achieves substantially better user satisfaction scores of 3.62 compared to CodeAct agent’s 2.57 (+41% improvement; automatically measured through user simulators that evaluate preference alignment, communication, etc.). The results highlight the importance of modeling user mental state in real-world software development, especially when user instructions are high-level and underspecified and preferences and constraints are not stated explicitly.
@@
Finally, we conduct a three-week in-the-wild observational human study with 17 professional developers using a ToM-enhanced OpenHands CLI (Wang et al., 2025) on their everyday, self-chosen coding tasks (rather than predefined tasks as in prior work (Wu et al., 2025; Qian et al., 2025a)). Our primary goal is to evaluate the interaction-level utility of ToM suggestions in professional workflows: each time the system produces a ToM suggestion, developers mark it as Accept, Almost right, let me modify it, or Reject. Across 209 sessions (174 ToM suggestions), developers accept or partially accept ToM suggestions 86% of the time. Besides the high acceptance rate, we learn from the developers’ feedback that ToM agent can often provide useful and valuable suggestions that make users workflow more efficient (e.g., “Please add pytest for the new function” or “keep your code edit minimal”). Furthermore, ToM agent can even provide novel and preference aligned suggestions (e.g., “refactor the code following Linux development philosophy”). Our work sheds light on building more proactive and personalized software engineering agents that can adapt to different users and tasks.
2 Related Work
Software Engineering Agents
Large language models enable competitive AI agents for software engineering. Systems like SWE-agent (Yang et al., 2024), CodeAct (Wang et al., 2024), demonstrate impressive automation capabilities through agent-computer interfaces and executable code actions (Jimenez et al., 2024). The standard benchmark for software engineering agents is SWE-bench (Jimenez et al., 2024), which evaluates agents’ ability to handle well-specified software engineering tasks, and use unit tests to decide the task success. However, these systems only interact with environments and overlook human developer interaction. Recently, ClarifyGPT (Mu et al., 2023) addresses requirement ambiguity through two-step consistency checks but focuses only on simple function generation. Vijayvargiya et al. (2026) introduce Ambiguous SWE-bench, showing that interaction helps agents handle underspecified software engineering tasks. However, as the benchmark operationalizes ambiguity primarily by underspecifying the current instruction, it does not capture the long-term memory and preference-tracking challenges that arise in multi-session settings.
Theory of Mind and Personalization in AI Systems
Theory of mind (ToM) is crucial for AI systems engaging with humans. Li et al. (2023) show that LLM-based agents with ToM capabilities better coordinate in multi-agent collaboration, essential for complex software development (Qian et al., 2024). While popular benchmarks like SOTOPIA (Zhou et al., 2024) evaluate social intelligence and FANToM (Kim et al., 2023) stress-test machine ToM, the ToM abilities of software engineering agents remain underexplored despite requiring close human-agent collaboration. Similarly, many personalization techniques have been proposed, including personalized reinforcement learning (Li et al., 2024), parameter-efficient conversation personalization (Magister* et al., 2025), user embeddings (Ning et al., 2024). However, how to effectively build personalized SWE agents is still an open question.
Agent Memory Systems
Effective memory management is critical for long-term AI agent interactions. RAG systems suffer from context pollution and fail to capture nuanced information (Letta, 2024). Recent advances include MemGPT (Packer et al., 2024) with dual-tier memory hierarchies, Mem0 (Chhikara et al., 2025) achieving faster context retrieval through dynamic consolidation, and MemoRAG (Qian et al., 2025b) addressing context pollution via draft generation. However, existing systems focus on general conversation contexts, leaving a gap in memory architectures for software engineering where agents must maintain complex mental models of user preferences, coding styles, and evolving requirements across sessions.
@@ 3 ToM-SWE: Pairing SWE Agent with ToM Agent
Consider existing setups of software engineering agents such as SWE-agent (Yang et al., 2024) and OpenHands CodeAct (Wang et al., 2024, 2025). At time step $t$ in coding session $i$, an agent receives an observation $o_{t}\in O$ either from the environment (e.g., terminal output, file contents, test results) or the user (e.g., user instructions, feedback). Then the agent takes an action $a_{t}\in A$ (e.g., code edits, shell commands) following some policy $\pi(a_{t}|c^{i}_{t})$, where $c^{i}_{t}=(o_{1},a_{1},\ldots,o_{t-1},a_{t-1},o_{t})$ is the context available to the agent in the coding session $i$ until time step $t$.
@@
Figure 3: Overview of the stateful SWE benchmark. We first collect profiles following the three steps outlined above. We then create instances by pairing the user simulator of different profiles with SWE-bench issues. Original SWE-bench issue descriptions are converted into vague user instructions (similar to Ambiguous SWE-bench (Vijayvargiya et al., 2026)), but our user simulator conditions on personalized profiles with distinct interaction styles and coding preferences, whereas Ambiguous SWE-bench uses fixed user simulator prompts. The agent solves the task under the same environment and tests of the original instance with access to the previous interaction histories with the user and could interact with the simulated user for extra information.
@@ 5.1 Experiment Setup
Our experiments involve three agent variants: (1) CodeAct agent: The baseline implementation following Wang et al. (2024), which operates without explicit user modeling capabilities. (2) RAGCodeAct agent: An enhanced version that encourages the coding agent to proactively retrieve relevant information from previous interaction history, serving as the single agent paradigm to manage both the coding task and user modeling task simultaneously. (3) TomCodeAct agent: Our proposed approach that pair coding agent with a theory-of-mind agent, maintaining explicit user mental models and adapting behavior accordingly. We evaluate all agents using three state-of-the-art language models: Claude Sonnet 4, Claude 3.7 Sonnet and Qwen3-480B (Qwen3). All agents interact with the same simulated users and are prompted to ask for clarifications when uncertain. We use Claude Sonnet 4 for the theory-of-mind agent across all experiments and use BM25 to retrive relevant information for both the RAGCodeAct agent and the TomCodeAct agent (see Appendix A.8 for complete model specifications and hyperparameters).
Figure 4: Performance comparison based on task resolved rates (measured by passed unit tests provided in the SWE-bench issues). TomCodeAct agent consistently outperforms CodeAct agent across both benchmarks and model variants, with the largest performance gap observed in the Stateful SWE benchmark using Claude Sonnet 4.
@@
As described in Section 4, we evaluate our approach on two complementary benchmarks: the Ambiguous SWE benchmark and our Stateful SWE benchmark. Both benchmarks use GPT-5 powered simulated users and evaluate agents over 500 instances with a maximum of 100 interaction turns per task. Note that the user simulator has access to both the complete issue description and the hints for the issue from the original SWE-bench issues. Together, these benchmarks evaluate complementary scenarios: ambiguous formal specifications requiring clarification versus informal user instructions with available context. We report task resolved rates (i.e., the percentage of tasks passed the unit tests provided in the SWE-bench issues) for both benchmarks and additionally report user simulator satisfaction scores for the Stateful SWE benchmark (i.e., the satisfaction scores generated by the LLM user simulator)  [^1] Please refer to the Appendix A.10 for more details on the scoring process of user simulators..
@@ 5.2 Benchmark Results
Figure 4 shows the resolved rates across all model-agent combinations. For example, TomCodeAct agent maintains its lead with 63.4% resolution rate using Claude Sonnet 4 versus CodeAct agent’s 51.9% on the Ambiguous SWE benchmark. And TomCodeAct agent achieves 59.7% resolution rate with Claude Sonnet 4 compared to CodeAct agent’s 18.1%, representing a 41.6 percentage point improvement on the Stateful SWE benchmark.
@@
Table 1 shows user satisfaction scores for the Stateful SWE benchmark. TomCodeAct agent achieves the highest satisfaction ratings across all models  [^2] Note that we tested a single-agent baseline that manages both coding tasks and ToM-related contexts. In preliminary experiments (N=100) with Claude Sonnet 4, this setup achieved only a 27.17% resolution rate on the stateful SWE bench. We attribute this underperformance to the difficulty of managing high-complexity contexts within a single agent and did not extend the baseline further due to budget constraints..
@@
Besides the success of the TomCodeAct agent, we additionally find that Theory-of-mind reasoning over past user-agent interactions is essential for understanding current user intent while simply retrieving raw session data provides limited help. We observe that while RAGCodeAct agent also has access to the previous raw session data, it still underperforms the ToMCodeAct agent across all models. This is especially true when the base model for SWE agent is not powerful enough to handle both the coding task and the user modeling task simultaneously. With Claude 3.7 Sonnet, RAGCodeAct agent even hurts the performance of the SWE agent (task resolved rate drops from 18.7% to 14.4%).
Relationship between Task Resolution and User Satisfaction
We conduct regression and correlation analyses to understand the relationship between task resolution rate and user satisfaction. We find a strong positive association: for the Claude 4 ToMCodeAct agent, the Pearson correlation coefficient between resolution and satisfaction is $r=0.74$ ($p<0.001$), and task resolution explains 55% of the variance in satisfaction scores ($R^{2}=0.55$). However, we also observe cases where the agent fails to resolve the task but user satisfaction is high (F+H) [^3] We also observe cases where tasks are resolved but satisfaction is only medium (S+M) due to preference violations or communication style (e.g., “Ignored the preferred typing style… Did not provide a descriptive commit message” and “You didn’t provide a brief summary…”). However, we find these cases happen less frequently (3% of the time), and there’s no statistically significant difference of the S+M rates among different agent-model configurations.. Specifically, we bucket satisfaction scores into High (3.5–5), Medium (2–3.5), and Low (1–2), and focus on this key disagreement case (F+H). Table 2 reports F+H rates across agent–model configurations. We find that the ToMCodeAct agent consistently achieves the highest F+H rates, suggesting a “good-effort bonus” where users reward strong process quality and preference alignment even without a final fix. Qualitatively, in F+H cases the simulator often credits meaningful progress and good communication (e.g., “precise fix with minimal back-and-forth, aligned well with my preferences” and “Asked for all key details up front (version, minimal repro, affected APIs), matching the preferred workflow”). To understand F+H cases in more detail, we manually examine 44 instances where the Claude 4 ToMCodeAct agent received high satisfaction despite failing the task; users were rewarding process quality rather than being misled about outcomes. Specifically, users valued (a) systematic, well-targeted investigation with clear technical reasoning (61% of F+H cases) and (b) clear communication and visible progress toward a solution (39% of F+H cases).
Table 2: Task resolution and user satisfaction disagreement rates (in %). F+H: failed to resolve the task but user satisfaction is high (higher is better).
CodeAct ToMCodeAct RAGCodeAct
[…]
Claude 3.7 3.7 19.3 5.0
Claude 4 8.8 21.5 14.0
Qwen3 5.9 15.4 8.0
@@ 5.3 Cost Efficiency of ToM Agent
For each ToM base model (GPT-5 nano, GPT-5 mini, GPT-5, Claude 3.7, Claude 4), we run the TomCodeAct agent on 100 sampled Stateful SWE-bench instances. We then report the resolved rate (%) alongside the average session cost (USD), which includes the full end-to-end cost of using the ToM agent during problem solving. As shown in Figure 5, even very efficient ToM base models substantially improve performance while adding only a small fraction of the overall session cost. For instance, GPT-5 nano reaches 38.0% at only \$0.02 per session. The top axis reports ToM overhead as a fraction of the average SWE cost per session (\$1.08); for example, using Claude 4 as the ToM base adds \$0.17 per session, i.e., about 16% overhead.
6 In-the-wild Human Study
To assess the real-world utility of TomCodeActAgent, we conducted a three-week observational field study with 17 software developers. Unlike a controlled laboratory experiment, this “in-the-wild” deployment allowed participants to interact with the agent within their familiar CLI environment while working on their own unscripted, day-to-day coding tasks.
The primary objective was to evaluate the interaction-level utility of the ToM agent’s suggestions in a professional context. Given the diverse and private nature of the participants’ work, we prioritized an observational design over a comparative control-group study to maximize ecological validity. Consequently, we focus our analysis on user-acceptance metrics, specifically, whether developers accepted, modified, or rejected the ToM agent’s suggestions. Here we don’t broad causal claims about productivity improvements.
In total, we recorded 209 coding sessions. Within these sessions, the ToM agent proactively provided 174 suggestions based on the inferred user state. Due to strict privacy constraints, we did not collect codebase content or specific dialogue, focusing exclusively on the developers’ behavioral responses to the agent’s interventions.
Figure 5: Resolved rate vs avg. ToM cost per session in stateful SWE benchmark. The orange line highlights the Pareto frontier: configurations that are not dominated by another option with both lower cost and higher resolved rate. The top axis shows ToM overhead as a percentage of the typical SWE session cost (\$1.08) per session.